You are here: Foswiki>PartnerObserving Web>ItalyQueue2012>ObsInaf20121208 (28 Feb 2013, EleanoraSani)Edit Attach

INAF LBC+MODS+LUCI queue observing, 2012 December 7-8

Observer: E. Sani, F. Cusano, M. Fumana
Telescope Operator: G. Becchetti
Support Astronomer: M. Edwars
Instrument: LBC/MODS1

Summary

Clear sky and no wind at the beginning of the night.
After 7.00 UT the humidity significantly increases reaching 87%.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Giallongo_27A MS0015.9+16 g-sloan, r-sloan 1h30m 4 OBs executed, we considered not valid the first OB due to elongated stars. COMPLETE
Standard SA92 g-sloan, r-sloan   airmass 1.3
Giallongo_27C MAC0744+39 g-sloan, r-sloan 3h COMPLETE
Standard SA98 g-sloan, r-sloan   airmass 1.2
Fritz_04 LensGG243 dual_grating 45m STD

00.25 Sky flat
  • The camera was not connected so we do not have WCS in the header of the images and we were no tracking
  • Tomorrow we are going to reacquire the sky flat check flat of this night

01.16 We slew to Giallongo_27A and run MS0015_9+16_GR_POINTING.ob
  • lbcrangeball
  • dofpia, /first, /X2
  • dopfia is not converging, we try again

01.33 Run MS0015_9+16_GR_POINTING.ob
  • lbcrangeball
  • dofpia, /first, /X2

01.38 We run MS0015_9+16_GR_a.ob
  • The 3rd image has elongated stars in the blue channel. We monitor the quality of incoming images.
  • images ==
  • In the blue images we see 'drop' features close to the brightest star
  • There was a running script generating problems with the telescope in binocular mode

2.17 Problem solved, the images are good

02.20 dopfia

02.40 Run MS0015_9+16_GR_d.ob *seeing on images ~0.8

03.21 dopfia

03.28 Run MS0015_9+16_GR_e.ob *seeing on images ~0.8-9

04.08 dopfia

04.14 Run MS0015_9+16_GR_f.ob

04.56 Run the acquisition of the standard SA92_gr.ob

05.01 We slew to Giallongo_27C and run MAC0744+39_GR_POINTING.ob *The telescope is not pointing, problem with the script, Preset LBT and Rotator Tracking were not selected also the Class Type Observations was setted on focus, so we had to create a new script

05.11 Launched correct MAC0744+39_GR_POINTING.ob *dopfia *lbcrangeball

05:20 Run MAC0744+39_GR_a.ob

06.00 dofpia

06.04 Run MAC0744+39_GR_b.ob

06.45 dofpia

06.51 Run MAC0744+39_GR_c.ob
  • seeing is 1" in both channels

07.30 dofpia

07.35 Run MAC0744+39_GR_d.ob

08.27 dofpia

08.30 Run MAC0744+39_GR_e.ob
  • seein is ~0.9" in both channels

09.13 dofpia

09.19 Run MAC0744+39_GR_f.ob
  • The Moon (illumination 32%) is rising. We monitor the background counts. In the last two frames there is an increase of the background of 10% in the blue channel and no significant differences for the red channel.
  • Images look ok, but lbc guider monitor show a strange issue (we have a screanshoot)

10.05 Slew to the standard SA98
  • dofpia
  • Run SA98_g_r.ob

10.23 SWITCH TO MODS

10:47 Even if MODS is already configured we can not start observations because of the shield camera temperature of LUCI. E. Solheid is checking the instrument.

11.20 We are authorized to use MODS and slew to Fritz_04A

11.30 acqMODS G191B2b.acq mods1r.20121208.0003
  • the STD script is wrong itis configured like a normal LS observation rather than that of a 5" slits STD. Thus running modsAlign -l we do not see the star in the field

11.38 acqMODS G191B2b.acq
  • modsAlign -l mods1r.20121208.0005.fits mods1r.20121208.0006.fits

11.50 execMods G191B2b.obs mods1r 0009 / mods1b 0003

12.02 acqMODS FritzLensGG243.acq

01.06 We stop the integration due to twilight

01.20 Dome closed
  • We can not perform MODS calibration because the LBTO crew has to monitor LUCI.

-- VincenzoTesta - 08 Dec 2012
Topic revision: r8 - 28 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback